# JWST Insights into Narrow-line Little Red Dots

### v1

@@ 5.2 Velocity Dispersion and Host Dynamical Mass
The tight correlation between black hole mass and central (stellar) velocity dispersion has been extensively examined and firmly established in the local Universe (e.g., J. Kormendy & L. C. Ho, 2013). Spatially resolved observations leveraging gravitational lensing have extended investigations of this relationship to (A. B. Newman et al., 2025). At even higher redshifts, measurement of often uses spatially integrated gas emission lines as proxies due to the limit of spatial resolution (e.g., R. Maiolino et al., 2024). These high-redshift studies suggest that this relation is fundamental and may have little evolution. Assuming that narrow-line LRDs are AGNs hosting low-mass SMBHs, we investigate whether they follow these two relations.
@@
Four of the narrow-line LRDs (except UID-16321) have medium-resolution spectral coverage of the [O iii] line, allowing us to measure the velocity dispersion of the ionizing gas (). As shown in Figure 6(b), is slightly larger than for the narrow-line LRDs, possibly due to the contribution from the central AGN. Therefore, we measure using the line. The spectral resolution of the medium-grating for compact sources at the spectral range of interest is , which is comparable to the derived intrinsic velocity dispersion. Correcting for the LSF introduces a significant source of uncertainty, which dominates the final error budget. In reality, the integrated measured using the emission line is not necessarily the same as the stellar velocity dispersion used in the local scaling relations. As done by H. Übler et al. (2023) and R. Maiolino et al. (2024), we use the empirical relation between observed and of low-redshift galaxies from R. Bezanson et al. (2018) to correct the and obtain a close estimate of . However, the – relation in R. Bezanson et al. (2018) is derived using massive galaxies, which may not be applicable to AGNs. Therefore, we also try to use the relation calibrated by L. C. Ho (2009, Equation 3) based on AGNs. These relations have a scatter of dex that is not considered in our error budget.
@@
Using the stellar velocity dispersion above and the morphological information from the PSF+Sérsic fitting in Section 3.3, we estimate the dynamical masses of the narrow-line LRD host galaxies. Only UID-8219 and UID-12329 have significant extended emission detected, while the other two sources can be well fitted by a single PSF model. Therefore, the PSF+Sérsic fitting result of UID-20547 and UID-22015 could have large uncertainty. Following H. Übler et al. (2023) and R. Maiolino et al. (2024) we adopt the equation:
@@
where and is the Sérsic index (M. Cappellari et al., 2006), and is the axis ratio (A. van der Wel et al., 2022), and is the effective radius. We use the , , and of the Sérsic component in the F150W or F200W band.
@@
The estimated and are reported in Table 2. The relations are shown in Figure 15 and compared with the high-redshift AGN samples from R. Maiolino et al. (2024) and the local relation from J. Kormendy & L. C. Ho (2013) and J. E. Greene et al. (2020). For both the – and – relations, the black holes appear undermassive relative to the local relation when adopting the R. Bezanson et al. (2018) correction. Adopting the L. C. Ho (2009) correction results in larger correction and therefore lower and values. This may be because, in the AGN case, energy injected by AGN feedback accelerates the NLR gas to velocities exceeding the stellar virial velocities (L. C. Ho, 2009). Under this correction, the narrow-line LRDs could either be consistent with the local – and – relations, or they may indeed host undermassive black holes. For UID-12329 and UID-20547, their stellar masses from the SED fitting assuming all emission is from galaxies are larger than the derived dynamical masses here (see Table 2), suggesting that the stellar masses are overestimated or the dynamical masses are underestimated. The black hole masses here are conservative upper limits. Therefore, in the AGN scenario, these narrow-line LRDs may represent a distinct population of AGNs that fall below the local – and – relations, potentially indicating deviations from the idea that the – relation is fundamental and already in place at high redshift. We notice that the current measurements suffer from large uncertainties. The expected for black holes with masses of is . Such a lower value can only be measured accurately by future higher resolution spectra.
5.3 Implications of Narrow-line LRDs
@@
Conversely, if narrow-line LRDs are AGNs, their low black hole masses () place them at the very low-mass end of the AGN population. These objects may represent an initial stage of the black hole growth, where accretion is super-Eddington but the black hole has yet to accumulate significant mass. The low fraction of narrow-line LRDs indicates that this evolutionary stage is short and the black hole grow very fast. As time goes on, their black holes may grow faster than the host galaxies, becoming more overmassive systems, resembling typical JWST AGNs (e.g., Y. Harikane et al., 2023; R. Maiolino et al., 2024; C.-H. Chen et al., 2025b).
Narrow-line LRDs may also represent the Type 2 counterparts of the broad-line LRDs, with the difference arising from torus obscuration. However, the striking similarity in SEDs between narrow-line and broad-line LRDs poses a puzzle. In local AGN populations, Type 1 and Type 2 sources typically exhibit distinct SEDs due to dust attenuation in the torus. If narrow-line LRDs are the Type 2 counterparts of broad-line LRDs, their SEDs should differ significantly. The observed similarity instead suggests either (1) a lack of a standard obscuring torus in LRDs (e.g., Z. Li et al., 2025), or (2) a scenario where both AGN and host galaxy contribute similarly to the V-shape continuum, decreasing orientation-dependent differences. The low fraction of narrow-line LRDs suggests that the broad-line region (BLR) of LRDs have a large covering factor (as suggested by R. Maiolino et al., 2025) or their torus have a small opening angle. Future investigation of the near-IR and mid-IR properties of narrow-line LRDs will be essential to test these hypotheses. It is also worth noting that some Type 2 AGNs in the local universe intrinsically lack a BLR, as their X-ray spectra reveal no intrinsic absorption (e.g., L. C. Ho et al., 2012; E. Pons & M. G. Watson, 2016). This scenario may potentially explain the similarity between the SED of narrow-line and broad-line LRDs, although it is difficult to confirm due to the lack of X-ray detections.
@@
5.4 Relation between Broad H Line and V-shape Continuum
@@
Spectroscopic follow-up of photometrically identified LRDs reveal that of them exhibit broad Balmer lines (e.g., L. J. Furtak et al., 2023; M. Killi et al., 2024; J. E. Greene et al., 2024; J. Matthee et al., 2024; B. Wang et al., 2024). On the other hand, among AGNs with broad lines, only show a LRD-like V-shape SED (e.g., A. J. Taylor et al., 2024; K. N. Hainline et al., 2025; I. Juodžbalis et al., 2025). We further investigate the connection between the broad line emission and V-shape continuum. Specifically, we examine the distribution and fraction of sources with broad H lines in the – plane.
@@
We use all sources with NIRSpec prism observation and high- or medium-resolution grating spectra covering H. The prism spectra enable us to eliminate the emission line contribution when measure the continuum slope. We also require the sources to be compact, satisfying the same criterion described in Section 3.1. In total, there are 491 sources. These sources are classified into three categories: (1) AGNs with obvious broad H line detections, defined by ; (2) sources without obvious broad H line detections with + 1; and (3) the remaining sources labeled as ‘vague’. We visually inspect the fitting result to ensure that the broad H line detections are real.
@@
The versus distribution of these sources is shown in Figure 16 (a). We divide the plane into 0.50.5 cells and calculate the fraction of broad-line AGNs in each grid cell, where vague sources are counted as broad-line AGNs with a weight of 0.5. The resultant fraction map is smoothed using a 2D Gaussian filter with a standard deviation of 1. The distribution of the broad H AGNs is similar to the result in K. N. Hainline et al. (2025) , spanning a wide range of . Compared to K. N. Hainline et al. (2025) who compiled AGNs from various studies, our sample is selected in a uniform manner, and we also subtract the contribution of emission lines to accurately measure the intrinsic continuum slope. Among the 55 broad-line AGNs, 24 () of them exhibit V-shape continuum with and . This fraction is slightly higher than that reported by previous studies (e.g., A. J. Taylor et al., 2024; K. N. Hainline et al., 2025; I. Juodžbalis et al., 2025), which may be due to the complex selection effects associated with the NIRSpec observations. The distribution of broad-line AGNs in the – plane appears nearly uniform, suggesting that broad-line LRDs are likely AGNs located at the red end of the rest-frame optical slope distribution. In terms of the broad-line AGNs fraction, a clear demarcation appears at : on the redder side, the fraction of AGNs increases significantly and becomes dominant, while on the bluer side, the AGN fraction drops sharply due to the increasing number of star-forming galaxies. Therefore, selecting sources with red rest-frame optical colors can yield broad-line AGNs with high purity, which is in line with the result in J. E. Greene et al. (2024). Such a selection is incomplete, as it misses a substantial number of AGNs located on the bluer end.
@@
Another important question is the redshift evolution of the LRD fraction in AGNs, which would provide valuable insights into the growth and obscuration of black holes across cosmic time. In Figure 16 (b) we show the redshift distributions of the broad H lines emitters and broad-line LRDs, along with the corresponding non-LRD fractions in each redshift bin. The number of LRDs increases from high to intermediate redshift, then turns over and declines at , consistent with the results in D. D. Kocevski et al. (2024). The non-LRD fraction show a increasing trend toward low redshift, supporting the idea that LRDs represent an early phase of AGN evolution that is more prevalent at high redshift. This observed non-LRD fraction evolution is consistent with the predictions of K. Inayoshi (2025), who suggest that LRD correspond the very first phase of SMBH growth, and their fraction declines as SMBH undergone multiple accretion episodes over cosmic time.
6 Conclusion
@@
In summary, our study reveals that narrow-line LRDs represent a distinct population. They may either be dusty, compact star-forming galaxies or AGNs with very low-mass black holes. Their spectral and structural properties provide new insights into the physical nature and diversity of LRDs. An unbiased study of flux-limited sample of narrow-line LRDs using deep slitless surveys (e.g., COSMOS-3D, NEXUS, and SAPPHIRES; K. Kakiichi et al., 2024; Y. Shen et al., 2024; X. Lin et al., 2025; F. Sun et al., 2025; M.-Y. Zhuang et al., 2025) will reduce selection effects and provide a better constraint on their fraction to improve our understanding of their role within the LRD population. Future deep JWST spectroscopy with higher resolution and sensitivity, along with deeper UV imaging, will be crucial for better understanding the nature of these sources.
@@ Appendix A Balmer Decrement
Here we introduce the calculation of nebular extinction using Balmer Decrement. We derive the dust reddening following the extinction law of D. Calzetti et al. (2000):
where and are the reddening curves evaluated at the H and H wavelengths, respectively. The factor is the color excess defined for H and H. is the observed Balmer decrement and is the intrinsic Balmer decrement. We assume the value of , corresponding to a temperature K and an electron density for Case B recombination (D. E. Osterbrock, 1989). We adopt the values derived from the best-fit CIGALE model in Section 4.1.

### v2

@@ V.2 Velocity Dispersion and Host Dynamical Mass
The tight correlation between black hole mass and central (stellar) velocity dispersion has been extensively examined and firmly established in the local Universe (e.g., Kormendy2013). Spatially resolved observations leveraging gravitational lensing have extended investigations of this relationship to (2025arXiv250317478N). At even higher redshifts, measurement of often uses spatially integrated gas emission lines as proxies due to the limit of spatial resolution (e.g., 2024A&A...691A.145M). These high-redshift studies suggest that this relation is fundamental and may have little evolution. Assuming that narrow-line LRDs are AGNs hosting low-mass SMBHs, we investigate whether they follow these two relations.
@@
Four of the narrow-line LRDs (except UID-16321) have medium-resolution spectral coverage of the [O iii] line, allowing us to measure the velocity dispersion of the ionizing gas (). As shown in Figure 6(a), is slightly larger than for the narrow-line LRDs, possibly due to the contribution from the central AGN. Therefore, we measure using the line. The spectral resolution of the medium-grating for compact sources at the spectral range of interest is , which is comparable to the derived intrinsic velocity dispersion. Correcting for the LSF introduces a significant source of uncertainty, which dominates the final error budget. In reality, the integrated measured using the emission line is not necessarily the same as the stellar velocity dispersion used in the local scaling relations. As done by 2023A&A...677A.145U and 2024A&A...691A.145M, we use the empirical relation between observed and of low-redshift galaxies from 2018ApJ...868L..36B to correct the and obtain a close estimate of . However, the – relation in 2018ApJ...868L..36B is derived using massive galaxies, which may not be applicable to AGNs. Therefore, we also try to use the relation calibrated by 2009ApJ...699..638H based on AGNs. These relations have a scatter of dex that is not considered in our error budget.
@@
Using the stellar velocity dispersion above and the morphological information from the PSF+Sérsic fitting in Section III.3, we estimate the dynamical masses of the narrow-line LRD host galaxies. Only UID-8219 and UID-12329 have significant extended emission detected, while the other two sources can be well fitted by a single PSF model. Therefore, the PSF+Sérsic fitting result of UID-20547 and UID-22015 could have large uncertainty. Following 2023A&A...677A.145U and 2024A&A...691A.145M we adopt the equation:
@@
where and is the Sérsic index (2006MNRAS.366.1126C), and is the axis ratio (2022ApJ...936....9V), and is the effective radius. We use the , , and of the Sérsic component in the F150W or F200W band.
@@
The estimated and are reported in Table 2. The relations are shown in Figure 16 and compared with the high-redshift AGN samples from 2024A&A...691A.145M and the local relation from Kormendy2013 and Greene2020. For both the – and – relations, the black holes appear undermassive relative to the local relation when adopting the 2018ApJ...868L..36B correction. Adopting the 2009ApJ...699..638H correction results in larger correction and therefore lower and values. This may be because, in the AGN case, energy injected by AGN feedback accelerates the NLR gas to velocities exceeding the stellar virial velocities (2009ApJ...699..638H). Under this correction, the narrow-line LRDs could either be consistent with the local – and – relations, or they may indeed host undermassive black holes. For UID-12329 and UID-20547, their stellar masses from the SED fitting assuming all emission is from galaxies are larger than the derived dynamical masses here (see Table 2), suggesting that the stellar masses are overestimated or the dynamical masses are underestimated. The black hole masses here are conservative upper limits. Therefore, in the AGN scenario, these narrow-line LRDs may represent a distinct population of AGNs that fall below the local – and – relations, potentially indicating deviations from the idea that the – relation is fundamental and already in place at high redshift. We notice that the current measurements suffer from large uncertainties. The expected for black holes with masses of is . Such a lower value can only be measured accurately by future higher resolution spectra.
V.3 Implications of Narrow-line LRDs
@@
Conversely, if narrow-line LRDs are AGNs, their low black hole masses () place them at the very low-mass end of the AGN population. These objects may represent an initial stage of the black hole growth, where accretion is super-Eddington but the black hole has yet to accumulate significant mass. The low fraction of narrow-line LRDs indicates that this evolutionary stage is short and the black hole grow very fast. As time goes on, their black holes may grow faster than the host galaxies, becoming more overmassive systems, resembling typical JWST AGNs (e.g., 2023ApJ...959...39H; 2024A&A...691A.145M; 2025ApJ...983...60C).
Narrow-line LRDs may also represent the Type 2 counterparts of the broad-line LRDs, with the difference arising from torus obscuration. However, the striking similarity in SEDs between narrow-line and broad-line LRDs poses a puzzle. In local AGN populations, Type 1 and Type 2 sources typically exhibit distinct SEDs due to dust attenuation in the torus. If narrow-line LRDs are the Type 2 counterparts of broad-line LRDs, their SEDs should differ significantly. The observed similarity instead suggests either (1) a lack of a standard obscuring torus in LRDs (e.g., 2025ApJ...980...36L), or (2) a scenario where both AGN and host galaxy contribute similarly to the V-shape continuum, decreasing orientation-dependent differences. The low fraction of narrow-line LRDs suggests that the BLR of LRDs have a large covering factor (as suggested by 2025MNRAS.538.1921M) or their torus have a small opening angle. Future investigation of the near-IR and mid-IR properties of narrow-line LRDs will be essential to test these hypotheses. It is also worth noting that some Type 2 AGNs in the local universe intrinsically lack a BLR, as their X-ray spectra reveal no intrinsic absorption (e.g., 2012ApJ...759L..16H; 2016A&A...594A..72P). This scenario may potentially explain the similarity between the SED of narrow-line and broad-line LRDs, although it is difficult to confirm due to the lack of X-ray detections.
@@
V.4 Relation between Broad H Line and V-shape Continuum
@@
Spectroscopic follow-up of photometrically identified LRDs reveal that of them exhibit broad Balmer lines (e.g., 2023ApJ...952..142F; 2024A&A...691A..52K; 2024ApJ...964...39G; 2024ApJ...963..129M; 2024arXiv240302304W). On the other hand, among AGNs with broad lines, only show a LRD-like V-shape SED (e.g., 2024arXiv240906772T; 2025ApJ...979..138H; 2025arXiv250403551J). We further investigate the connection between the broad line emission and V-shape continuum. Specifically, we examine the distribution and fraction of sources with broad H lines in the – plane.
@@
We use all sources with NIRSpec prism observation and high- or medium-resolution grating spectra covering H. The prism spectra enable us to eliminate the emission line contribution when measure the continuum slope. We also require the sources to be compact, satisfying the same criterion described in Section III.1. In total, there are 491 sources. These sources are classified into three categories: (1) AGNs with obvious broad H line detections, defined by ; (2) sources without obvious broad H line detections with + 1; and (3) the remaining sources labeled as ‘vague’. We visually inspect the fitting result to ensure that the broad H line detections are real.
@@
The versus distribution of these sources is shown in Figure 17 (a). We divide the plane into 0.50.5 cells and calculate the fraction of broad-line AGNs in each grid cell, where vague sources are counted as broad-line AGNs with a weight of 0.5. The resultant fraction map is smoothed using a 2D Gaussian filter with a standard deviation of 1. The distribution of the broad H AGNs is similar to the result in 2025ApJ...979..138H , spanning a wide range of . Compared to 2025ApJ...979..138H who compiled AGNs from various studies, our sample is selected in a uniform manner, and we also subtract the contribution of emission lines to accurately measure the intrinsic continuum slope. Among the 55 broad-line AGNs, 24 () of them exhibit V-shape continuum with and . This fraction is slightly higher than that reported by previous studies (e.g., 2024arXiv240906772T; 2025ApJ...979..138H; 2025arXiv250403551J), which may be due to the complex selection effects associated with the NIRSpec observations. The distribution of broad-line AGNs in the – plane appears nearly uniform, suggesting that broad-line LRDs are likely AGNs located at the red end of the rest-frame optical slope distribution. In terms of the broad-line AGNs fraction, a clear demarcation appears at : on the redder side, the fraction of AGNs increases significantly and becomes dominant, while on the bluer side, the AGN fraction drops sharply due to the increasing number of star-forming galaxies. Therefore, selecting sources with red rest-frame optical colors can yield broad-line AGNs with high purity, which is in line with the result in 2024ApJ...964...39G. Such a selection is incomplete, as it misses a substantial number of AGNs located on the bluer end.
@@
Another important question is the redshift evolution of the LRD fraction in AGNs, which would provide valuable insights into the growth and obscuration of black holes across cosmic time. In Figure 17 (b) we show the redshift distributions of the broad H lines emitters and broad-line LRDs, along with the corresponding non-LRD fractions in each redshift bin. The number of LRDs increases from high to intermediate redshift, then turns over and declines at , consistent with the results in 2024arXiv240403576K. The non-LRD fraction show a increasing trend toward low redshift, supporting the idea that LRDs represent an early phase of AGN evolution that is more prevalent at high redshift. This observed non-LRD fraction evolution is consistent with the predictions of 2025arXiv250305537I, who suggest that LRD correspond the very first phase of SMBH growth, and their fraction declines as SMBH undergone multiple accretion episodes over cosmic time.
VI Conclusion
@@
In summary, our study reveals that narrow-line LRDs represent a distinct population. They may either be dusty, compact star-forming galaxies or AGNs with very low-mass black holes. Their spectral and structural properties provide new insights into the physical nature and diversity of LRDs. An unbiased study of flux-limited sample of narrow-line LRDs using deep slitless surveys (e.g., COSMOS-3D, NEXUS, and SAPPHIRES; 2024jwst.prop.5893K; 2024arXiv240812713S; 2025arXiv250408039L; 2025arXiv250315587S; 2025arXiv250520393Z) will reduce selection effects and provide a better constraint on their fraction to improve our understanding of their role within the LRD population. Future deep JWST spectroscopy with higher resolution and sensitivity, along with deeper UV imaging, will be crucial for better understanding the nature of these sources.
The data and code of this article are available on Zenodo under an open-source Creative Commons Attribution license: https://doi.org/10.5281/zenodo.18149018 (catalog doi:10.5281/zenodo.18149018). The GitHub repository URL is https://github.com/Zijian-astro/NL-LRD-2025-Data (catalog NL-LRD-2025-Data).
@@ Appendix A Balmer Decrement
Here we introduce the calculation of nebular extinction using Balmer Decrement. We derive the dust reddening following the extinction law of 2000ApJ...533..682C:
where and are the reddening curves evaluated at the H and H wavelengths, respectively. The factor is the color excess defined for H and H. is the observed Balmer decrement and is the intrinsic Balmer decrement. We assume the value of , corresponding to a temperature K and an electron density for Case B recombination (1989agna.book.....O). We adopt the values derived from the best-fit CIGALE model in Section IV.1.